# Accelerating Goal-Conditioned Reinforcement Learning Algorithms and Research

### v3

@@ 3.2 Policy Learning
We use a DDPG-style policy extraction loss to learn a goal-conditioned policy, $\pi_{\theta}(a|s,g)$, by optimizing the critic $f_{\phi,\psi}$ (Lillicrap et al.,, 2016):
@@ 4.1 JaxGCRL speedup on a single GPU
We compare the proposed fully JIT-compiled implementation of CRL in JaxGCRL to the original implementation from Eysenbach et al., (2022). In particular, Fig. 1 shows the performance in Ant environment along with the experiment’s wall-clock time. The speedup in this configuration is 22-fold, with the new implementation reaching a training speed of over $16500$ environment steps per second with a 1:16 update to data (UTD) ratio. For results with other UTD, refer to Section 5.6. Complete speedup results are provided in Section A.1. Importantly, BRAX physics simulator differs from the original MuJoCo, so performance numbers here may vary slightly from prior work.
@@ 4.2 JaxGCRL Environments in the Benchmark
Figure 2: JaxGCRL benchmark: New suite of GPU-accelerated environments for studying GCRL. In this setting, the agent does not receive any rewards or demonstrations, making some of these tasks an excellent testbed for studying exploration and long-horizon reasoning. Our accompanying implementation of GCRL algorithms trains with more than $15$K environment steps per second on a single GPU, enabling rapid experimentation.
To evaluate the performance of GCRL methods, we propose JaxGCRL benchmark consisting of 8 diverse continuous control environments. These environments range from simple, ideal for quick checks, to complex, requiring long-term planning and exploration for success. The following list provides a brief description of each environment, with the technical details summarized in Table 1:
Reacher (Brockman et al.,, 2022).
@@
Half Cheetah (Wawrzyński,, 2009).
@@
In Fig. 3, we report the baseline results for the proposed benchmark. It is worth noting that for most environments, experiments involving 50M environment steps can be completed in less than an hour.
@@ 4.3 Contrastive RL design choices
Measuring the similarity between samples can be achieved in various ways, resulting in potentially different agent behaviors. Our analysis in following sections include cosine similarity (Chen et al., 2020a, ), dot product  (Radford et al.,, 2021), and negative $L_{1}$ and $L_{2}$ distance (Hu et al.,, 2023), detailed list can be found in Section A.2. Even though there is no consensus on the choice of energy functions for temporal representations, recent works showed that they should abide by quasimetric properties (Wang et al.,, 2023; Myers et al., 2024b, ).
@@
Beside InfoNCE-type losses (van den Oord et al.,, 2019; Eysenbach et al.,, 2022), we evaluate FlatNCE-like losses (Chen et al.,, 2021), and a Monte Carlo version of Forward-Backward unsupervised loss (Touati and Ollivier,, 2021). Additionally, we test novel objectives inspired by preference optimization for large language models (Calandriello et al.,, 2024). Specifically, we evaluate DPO, IPO, and SPPO, which increase the scores of positive samples and reduce the scores of negative ones. A full list of contrastive objectives can be found in Section A.2.
@@
Scaling neural network architectures to improve performance is a common practice in other areas of deep learning, but it remains relatively underexplored in RL models, as noted in Nauman et al., 2024a ; Nauman et al., 2024b . Recently, Zheng et al., (2024) showed that CRL might benefit from deeper and wider architectures with Layer Normalization (Ba et al.,, 2016) for offline CRL in pixel-based environments; we want to examine whether this also holds for online state-based settings.
@@ 5 Examples of Fast Experiments Possible with the New Benchmark
The goal of our experiments is twofold: (1) to establish a baseline for the proposed JaxGCRL environments, and (2) to evaluate CRL performance in relation to key design choices. In Section 5.1, we define setup that is used for most of the experiments unless explicitly stated otherwise. First in Section 5.2, we report baseline results on JaxGCRL. Second, in Sections 5.3 and 5.4, we try to understand the influence of design choices on CRL learning performance. Third, in Section 5.5, we asses those design choices in a data-rich setting with 300M environment steps. Lastly, in Section 5.6, we explore the relation between performance and UTD.
@@ 5.1 Experimental Setup
Our experiments use JaxGCRL suite of simulated environments described in Section 4.2. We evaluate algorithms in an online setting for $50$M environment steps. We compare CRL with Soft Actor-Critic (SAC) (Haarnoja et al.,, 2018), SAC with Hindsight Experience Replay (HER) (Andrychowicz et al.,, 2017), TD3 (Fujimoto et al.,, 2018), TD3+HER, and PPO (Schulman et al.,, 2017). For algorithms with HER, we use final strategy relabelling, i.e. relabeling goals with states achieved at the end of the trajectory. In the majority of experiments, we use CRL with L2 energy function, symmetric InfoNCE objective, and a tuneable entropy coefficient for all methods. See Appendix B for details. We use two performance metrics: success rate and time near goal. Success rate measures whether the agent reached the goal at least once during the episode, while time near goal indicates how long the agent stayed close to it. We use a sparse reward for all baselines, with $r=1$ when the agent is in goal proximity and $r=0$ otherwise. We define goal-reaching as achieving proximity below the goal distance threshold defined in Table 1. The implementations of PPO and SAC are sourced from the Brax repository (Freeman et al.,, 2021), while TD3, HER, and CRL are partially based on Brax.
@@ 5.2 JaxGCRL benchmark results
We establish baseline results for JaxGCRL with all algorithms in Fig. 3. Clearly, CRL achieves the highest performance across tested methods, resulting in non-trivial policies even in the hardest tasks in terms of high-dimensional state and action spaces (Humanoid) and exploration (Pusher and Ant Push). However, the performance in these challenging environments is low, indicating room for improvement for future contrastive RL methods. As expected, HER can improve performance for both TD3 and SAC. In contrast, PPO performs poorly across all tasks, likely due to the challenges posed by the sparse reward setting. Additional experiments on JaxGCRL environments with design choices discussed in the following sections can be found in Section A.4.
Figure 3: Baseline results in JaxGCRL benchmark. Success rates of all the baseline algorithms for $50$M environment steps for every JaxGCRL environment. CRL outperforms other baselines in most of the environments. The training speed is a function of the environment complexity, method complexity, and physics backend; see Section A.4. Specifically, due to differences in how each method works, the speed varies greatly in the same environments; this can be best seen with the PPO method being significantly faster than others due to it not using a replay buffer, which frees up GPU memory for more parallel environment simulations. Results are reported as the interquartile mean (IQM) along with its standard error, based on 10 seeds.
@@ 5.3 Contrastive objectives and energy functions comparison
Contrastive objective and energy function are the two main components of contrastive methods, serving as the primary drivers of their final performance. We evaluate CRL with 10 different contrastive objectives, as defined in Section 4.3 across five environments: Ant Soccer, Ant, Ant Big Maze, Ant U-Maze, Ant Push, and Pusher, and report aggregated performance. For the energy function, we use L2, as it consistently resulted in the highest performance for the CRL method, especially regarding time near goal, see Section A.2. Additionally, we apply logsumexp regularization with a coefficient of 0.1 to each objective. This auxiliary objective is essential, as without it, the performance of InfoNCE deteriorates significantly (Eysenbach et al.,, 2022). The analysis presented in Fig. 4 reveals that the originally proposed NCE-binary objective, along with forward-backward, IPO, and SPPO, are the least effective objectives among those evaluated. However, for other InfoNCE-derived objectives, it is difficult to determine the best one, as their performance is similar. Interestingly, CRL seems fairly robust to the choice of contrastive objective.
@@ 5.4 Scaling the Architecture
In this section, we explore how increasing the size of actor and critic networks, in terms of both depth and width, influences CRL performance. We evaluate the aggregated performance during the final $10$M steps of a $50$ million-step training process across three environments: Ant, Ant Soccer, and Ant U-Maze. We use the L2, Symmetric InfoNCE and logsumexp regularizer coefficient of $0.1$ for all architecture sizes.
(a) Neurons per layer $=$ 256
(b) Neurons per layer $=$ 512
(c) Neurons per layer $=$ 1024
@@
We present the results of this scaling experiment in Fig. 6 and observe that increasing both the width and depth tends to increase performance. However, performance does not further improve when increasing the depth for width = $1024$ neurons. Our next experiment studies whether layer normalization can stabilize the performance of these biggest networks (width of $1024$ neurons, depth of $4$). Indeed, the results in Fig. 6 show that adding layer normalization before every activation allows better scaling properties, especially for bigger networks.
@@ 5.5 Scaling the data
We evaluate the benefits CRL gains from training in a data-rich setting. In particular, we report performance for large architectures (studied in Section 5.4) with different combinations of energy functions and contrastive objectives for $300$M environment steps in Fig. 8. We observe that the L2 energy function with InfoNCE objective configuration outperforms all others by a substantial margin, leading to a higher success rate and time near the goal across three locomotion tasks. Interestingly, the dot product energy function performs best in the object manipulation task (Ant Soccer). This indicates that only a subset of a wide array of design choices performs well when scaling CRL. Additionally, there is still room for improvement in scaling CRL with data, as the success rate in Ant Soccer and Ant Big Maze remains around $40\%$. For additional experiments, refer to Section A.6
@@ 5.6 Gradient updates to data ratio
JaxGCRL enables efficient execution of extensive experiments. Leveraging this capability, we explore the effect of the model’s update frequency in CRL by evaluating a range of UTD ratios. In particular, we examine ratios ($1$:$1$, $1$:$8$, $1$:$16$, $1$:$24$, $1$:$32$, $1$:$49$, $1$:$48$) in five environments: Ant, Ant Soccer, Ant U-Maze, Pusher Hard, and Ant Push. Interestingly, we only observe a significant increase in performance for Pusher Hard with a higher number of updates, while in other environments, it leads to decreased or similar performance. With a UTD ratio of $1$:$16$, our code is $\times2$ faster than prior implementations, and with a lower frequency of gradient updates, it can be even faster.
@@ A.1 Speedup comparison across various numbers of parallel environments
We present an extended version of the plot from Fig. 1, with additional experiments, as depicted on the left side of Fig. 9. Each experiment was run for 10M environment steps, and for the original repository, we varied the number of parallel actors for data collection, testing configurations with 4, 8, 16, and 32 actors, each running on separate CPU threads. Each configuration was tested with three different random seeds, and we present the results along with the corresponding standard deviations. The novel repository used 1024 actors for data collection. We used NVIDIA V100 GPU for this experiment.
A notable observation from these experiments is the variation in success rates associated with different numbers of parallel actors. We hypothesize that this discrepancy arises due to the increased diversity of data supplied to the replay buffer as the number of independent parallel environments increases, leading to more varied experiences for each policy update. We conducted similar experiments using our method with varying numbers of parallel environments to further investigate. The results are presented on the right side of Fig. 9. This observation, while interesting, is beyond the scope of the current work and is proposed as an area for further investigation.
@@ A.3 Energy functions results
The performance of contrastive learning is sensitive to energy function choice (Sohn,, 2016). This section aims to understand how different energy functions impact CRL performance. In particular, we evaluate five energy functions: L1, L2, L2 w/o sqrt, dot product and cosine. For every energy function, we use symmetric InfoNCE as a contrastive objective, with a 0.1 logsumexp penalty coefficient.
@@
In Fig. 10, we report the performance of every energy function for four different ant environments and pusher. We find that p-norms and dot-product significantly outperform cosine similarity. Additionally, removing the root square from the L2 norm (L2 w/o sqrt) results in performance degradation, especially regarding time near goal. This modification makes the energy function no longer abide by the triangle inequality, which, as pointed out by (Myers et al., 2024b, ), is desirable for temporal contrastive features. Results per environment are reported in Fig. 11.
Fig. 11 shows the results of different energy functions per environment, success rate and time near goal. Clearly, no single energy function performs well in all the tested environments, as, for instance, L1 and L2, which perform well in Ant environments, work poorly in Pusher. In addition, we observed high variability in every configuration performance, as indicated by relatively wide standard errors.
@@ A.4 Benchmark performance
In this section, report additional results for more advanced architectures on the benchmark environment. In particular, we report the results for architecture with 4 hidden layers of size 1024 and Layer Normalization in Fig. 12. Unsurprisingly, the performance is significantly better in most environments, particularly for Humanoid. This suggests that a larger architecture is needed to effectively manage the high-dimensional state and action spaces involved in this environment.
@@ Appendix B Technical details
JaxGCRL is a fast implementation of state-based self-supervised reinforcement learning algorithms and a new benchmark of GPU-accelerated environments. Our implementation leverages the power of GPU-accelerated simulators (BRAX and MuJoCo MJX) (Freeman et al.,, 2021; Todorov et al.,, 2012) to reduce the time required for data collection and training, allowing researchers to run extensive experiments in a fraction of the time previously needed. The bottleneck of former self-supervised RL implementations was twofold. Firstly, data collection was executed on many CPU threads, as a single thread was often used for a single actor. This reduced the number of possible parallel workers, as only high-compute servers could run hundreds of parallel actors. Secondly, the necessity of data migration between CPU (data collection) and GPU (training) posed additional overhead. These two problems are mitigated by fully JIT-compiled algorithm implementation and execution of all environments and replay buffer operations directly on the GPU.

### v4

@@ 3.2 Policy Learning
We use a DDPG-style policy extraction loss to learn a goal-conditioned policy, $\pi_{\theta}(a|s,g)$, by optimizing the critic $f_{\phi,\psi}$ (lillicrap2016continuousa):
@@ 4.1 JaxGCRL speedup on a single GPU
We compare the proposed fully JIT-compiled implementation of CRL in JaxGCRL to the original implementation from eysenbach2022contrastive. In particular, Fig.˜1 shows the performance in Ant environment along with the experiment’s wall-clock time. The speedup in this configuration is 22-fold, with the new implementation reaching a training speed of over $16500$ environment steps per second with a 1:16 update to data (UTD) ratio. For results with other UTD, refer to Section˜5.6. Complete speedup results are provided in Section˜A.1. Importantly, BRAX physics simulator differs from the original MuJoCo, so performance numbers here may vary slightly from prior work.
@@ 4.2 JaxGCRL Environments in the Benchmark
Figure 2: JaxGCRL benchmark: New suite of GPU-accelerated environments for studying GCRL. In this setting, the agent does not receive any rewards or demonstrations, making some of these tasks an excellent testbed for studying exploration and long-horizon reasoning. Our accompanying implementation of GCRL algorithms trains with more than 1515K environment steps per second on a single GPU, enabling rapid experimentation.
To evaluate the performance of GCRL methods, we propose JaxGCRL benchmark consisting of 8 diverse continuous control environments. These environments range from simple, ideal for quick checks, to complex, requiring long-term planning and exploration for success. The following list provides a brief description of each environment, with the technical details summarized in Table˜1:
Reacher (brockman2022openai).
@@
Half Cheetah (wawrzynskiCatLikeRobotRealTime2009).
@@
In Fig.˜3, we report the baseline results for the proposed benchmark. It is worth noting that for most environments, experiments involving 50M environment steps can be completed in less than an hour.
@@ 4.3 Contrastive RL design choices
Measuring the similarity between samples can be achieved in various ways, resulting in potentially different agent behaviors. Our analysis in following sections include cosine similarity (chenSimpleFrameworkContrastive2020), dot product  (radford2021learning), and negative $L_{1}$ and $L_{2}$ distance (hu2023your), detailed list can be found in Section˜A.2. Even though there is no consensus on the choice of energy functions for temporal representations, recent works showed that they should abide by quasimetric properties (wang2023optimal; myers2024learning).
@@
Beside InfoNCE-type losses (oord2019representation; eysenbach2022contrastive), we evaluate FlatNCE-like losses (chenSimplerFasterStronger2021), and a Monte Carlo version of Forward-Backward unsupervised loss (touati2021learning). Additionally, we test novel objectives inspired by preference optimization for large language models (calandriello2024humanalignmentlargelanguage). Specifically, we evaluate DPO, IPO, and SPPO, which increase the scores of positive samples and reduce the scores of negative ones. A full list of contrastive objectives can be found in Section˜A.2.
@@
Scaling neural network architectures to improve performance is a common practice in other areas of deep learning, but it remains relatively underexplored in RL models, as noted in nauman2024overestimation; nauman2024bigger. Recently, zheng2024stabilizing showed that CRL might benefit from deeper and wider architectures with Layer Normalization (ba2016layer) for offline CRL in pixel-based environments; we want to examine whether this also holds for online state-based settings.
@@ 5 Examples of Fast Experiments Possible with the New Benchmark
The goal of our experiments is twofold: (1) to establish a baseline for the proposed JaxGCRL environments, and (2) to evaluate CRL performance in relation to key design choices. In Section˜5.1, we define setup that is used for most of the experiments unless explicitly stated otherwise. First in Section˜5.2, we report baseline results on JaxGCRL. Second, in Sections˜5.3 and 5.4, we try to understand the influence of design choices on CRL learning performance. Third, in Section˜5.5, we asses those design choices in a data-rich setting with 300M environment steps. Lastly, in Section˜5.6, we explore the relation between performance and UTD.
@@ 5.1 Experimental Setup
Our experiments use JaxGCRL suite of simulated environments described in Section˜4.2. We evaluate algorithms in an online setting for 5050M environment steps. We compare CRL with Soft Actor-Critic (SAC) (haarnoja2018soft), SAC with Hindsight Experience Replay (HER) (andrychowicz2017hindsight), TD3 (fujimotoAddressingFunctionApproximation2018), TD3+HER, and PPO (schulman2017proximal). For algorithms with HER, we use final strategy relabelling, i.e. relabeling goals with states achieved at the end of the trajectory. In the majority of experiments, we use CRL with L2 energy function, symmetric InfoNCE objective, and a tuneable entropy coefficient for all methods. See Appendix˜B for details. We use two performance metrics: success rate and time near goal. Success rate measures whether the agent reached the goal at least once during the episode, while time near goal indicates how long the agent stayed close to it. We use a sparse reward for all baselines, with $r=1$ when the agent is in goal proximity and $r=0$ otherwise. We define goal-reaching as achieving proximity below the goal distance threshold defined in Table˜1. The implementations of PPO and SAC are sourced from the Brax repository (freeman2021brax), while TD3, HER, and CRL are partially based on Brax.
@@ 5.2 JaxGCRL benchmark results
We establish baseline results for JaxGCRL with all algorithms in Fig.˜3. Clearly, CRL achieves the highest performance across tested methods, resulting in non-trivial policies even in the hardest tasks in terms of high-dimensional state and action spaces (Humanoid) and exploration (Pusher and Ant Push). However, the performance in these challenging environments is low, indicating room for improvement for future contrastive RL methods. As expected, HER can improve performance for both TD3 and SAC. In contrast, PPO performs poorly across all tasks, likely due to the challenges posed by the sparse reward setting. Additional experiments on JaxGCRL environments with design choices discussed in the following sections can be found in Section˜A.4.
Figure 3: Baseline results in JaxGCRL benchmark. Success rates of all the baseline algorithms for 5050M environment steps for every JaxGCRL environment. CRL outperforms other baselines in most of the environments. The training speed is a function of the environment complexity, method complexity, and physics backend; see Section˜A.4. Specifically, due to differences in how each method works, the speed varies greatly in the same environments; this can be best seen with the PPO method being significantly faster than others due to it not using a replay buffer, which frees up GPU memory for more parallel environment simulations. Results are reported as the interquartile mean (IQM) along with its standard error, based on 10 seeds.
@@ 5.3 Contrastive objectives and energy functions comparison
Contrastive objective and energy function are the two main components of contrastive methods, serving as the primary drivers of their final performance. We evaluate CRL with 10 different contrastive objectives, as defined in Section˜4.3 across five environments: Ant Soccer, Ant, Ant Big Maze, Ant U-Maze, Ant Push, and Pusher, and report aggregated performance. For the energy function, we use L2, as it consistently resulted in the highest performance for the CRL method, especially regarding time near goal, see Section˜A.2. Additionally, we apply logsumexp regularization with a coefficient of 0.1 to each objective. This auxiliary objective is essential, as without it, the performance of InfoNCE deteriorates significantly (eysenbach2022contrastive). The analysis presented in Fig.˜4 reveals that the originally proposed NCE-binary objective, along with forward-backward, IPO, and SPPO, are the least effective objectives among those evaluated. However, for other InfoNCE-derived objectives, it is difficult to determine the best one, as their performance is similar. Interestingly, CRL seems fairly robust to the choice of contrastive objective.
@@ 5.4 Scaling the Architecture
In this section, we explore how increasing the size of actor and critic networks, in terms of both depth and width, influences CRL performance. We evaluate the aggregated performance during the final 1010M steps of a $50$ million-step training process across three environments: Ant, Ant Soccer, and Ant U-Maze. We use the L2, Symmetric InfoNCE and logsumexp regularizer coefficient of $0.1$ for all architecture sizes.
(a) Neurons per layer == 256
(b) Neurons per layer == 512
(c) Neurons per layer == 1024
@@
We present the results of this scaling experiment in Fig.˜6 and observe that increasing both the width and depth tends to increase performance. However, performance does not further improve when increasing the depth for width = $1024$ neurons. Our next experiment studies whether layer normalization can stabilize the performance of these biggest networks (width of $1024$ neurons, depth of $4$). Indeed, the results in Fig.˜6 show that adding layer normalization before every activation allows better scaling properties, especially for bigger networks.
@@ 5.5 Scaling the data
We evaluate the benefits CRL gains from training in a data-rich setting. In particular, we report performance for large architectures (studied in Section˜5.4) with different combinations of energy functions and contrastive objectives for 300300M environment steps in Fig.˜8. We observe that the L2 energy function with InfoNCE objective configuration outperforms all others by a substantial margin, leading to a higher success rate and time near the goal across three locomotion tasks. Interestingly, the dot product energy function performs best in the object manipulation task (Ant Soccer). This indicates that only a subset of a wide array of design choices performs well when scaling CRL. Additionally, there is still room for improvement in scaling CRL with data, as the success rate in Ant Soccer and Ant Big Maze remains around $40\%$. For additional experiments, refer to Section˜A.6
@@ 5.6 Gradient updates to data ratio
JaxGCRL enables efficient execution of extensive experiments. Leveraging this capability, we explore the effect of the model’s update frequency in CRL by evaluating a range of UTD ratios. In particular, we examine ratios (11:11, 11:88, 11:$16$, 11:$24$, 11:$32$, 11:$49$, 11:$48$) in five environments: Ant, Ant Soccer, Ant U-Maze, Pusher Hard, and Ant Push. Interestingly, we only observe a significant increase in performance for Pusher Hard with a higher number of updates, while in other environments, it leads to decreased or similar performance. With a UTD ratio of 11:$16$, our code is $22\times$ faster than prior implementations, and with a lower frequency of gradient updates, it can be even faster.
@@ A.1 Speedup comparison across various numbers of parallel environments
We present an extended version of the plot from Fig.˜1, with additional experiments, as depicted on the left side of Fig.˜9. Each experiment was run for 10M environment steps, and for the original repository, we varied the number of parallel actors for data collection, testing configurations with 4, 8, 16, and 32 actors, each running on separate CPU threads. Each configuration was tested with three different random seeds, and we present the results along with the corresponding standard deviations. The novel repository used 1024 actors for data collection. We used NVIDIA V100 GPU for this experiment.
A notable observation from these experiments is the variation in success rates associated with different numbers of parallel actors. We hypothesize that this discrepancy arises due to the increased diversity of data supplied to the replay buffer as the number of independent parallel environments increases, leading to more varied experiences for each policy update. We conducted similar experiments using our method with varying numbers of parallel environments to further investigate. The results are presented on the right side of Fig.˜9. This observation, while interesting, is beyond the scope of the current work and is proposed as an area for further investigation.
@@ A.3 Energy functions results
The performance of contrastive learning is sensitive to energy function choice (sohn2016improved). This section aims to understand how different energy functions impact CRL performance. In particular, we evaluate five energy functions: L1, L2, L2 w/o sqrt, dot product and cosine. For every energy function, we use symmetric InfoNCE as a contrastive objective, with a 0.1 logsumexp penalty coefficient.
@@
In Fig.˜10, we report the performance of every energy function for four different ant environments and pusher. We find that p-norms and dot-product significantly outperform cosine similarity. Additionally, removing the root square from the L2 norm (L2 w/o sqrt) results in performance degradation, especially regarding time near goal. This modification makes the energy function no longer abide by the triangle inequality, which, as pointed out by (myers2024learning), is desirable for temporal contrastive features. Results per environment are reported in Fig.˜11.
Fig.˜11 shows the results of different energy functions per environment, success rate and time near goal. Clearly, no single energy function performs well in all the tested environments, as, for instance, L1 and L2, which perform well in Ant environments, work poorly in Pusher. In addition, we observed high variability in every configuration performance, as indicated by relatively wide standard errors.
@@ A.4 Benchmark performance
In this section, report additional results for more advanced architectures on the benchmark environment. In particular, we report the results for architecture with 4 hidden layers of size 1024 and Layer Normalization in Fig.˜12. Unsurprisingly, the performance is significantly better in most environments, particularly for Humanoid. This suggests that a larger architecture is needed to effectively manage the high-dimensional state and action spaces involved in this environment.
@@ Appendix B Technical details
JaxGCRL is a fast implementation of state-based self-supervised reinforcement learning algorithms and a new benchmark of GPU-accelerated environments. Our implementation leverages the power of GPU-accelerated simulators (BRAX and MuJoCo MJX) (freeman2021brax; todorov2012mujoco) to reduce the time required for data collection and training, allowing researchers to run extensive experiments in a fraction of the time previously needed. The bottleneck of former self-supervised RL implementations was twofold. Firstly, data collection was executed on many CPU threads, as a single thread was often used for a single actor. This reduced the number of possible parallel workers, as only high-compute servers could run hundreds of parallel actors. Secondly, the necessity of data migration between CPU (data collection) and GPU (training) posed additional overhead. These two problems are mitigated by fully JIT-compiled algorithm implementation and execution of all environments and replay buffer operations directly on the GPU.